OpenStudy (anonymous):

Suppose \[-\infty \le a < c < b \le∞\] and f:(a,b)rightarrow Real is continuous on (a,b). (a) If f is uniformly continuous on both (a,c) and (c,b), prove that f is uniformly continuous on (a,b). (b) Give an example to show that the conclusion in part (a) may be false if f is not continuous on (a,b).

OpenStudy (anonymous):

First the counter example f(x) =1 if x is in the interval [0, 1{ f(x) = -1 if x is in the interval [1, 2] Can you show that this f is a counter example?

OpenStudy (anonymous):

Let \[ \epsilon > 0 \text { and choose } \delta > 0 \text { so that } \\ | x - c| < \delta \text { implies } \, |f(x) - f(c) \le \frac {\epsilon}2 \\ | x - y| < \delta \text { implies } \, |f(x) - f(y) \le \frac {\epsilon}2 \\ \text { if both } x,y \text { are } in (a, c)\\ | x - y| < \delta \text { implies } \, |f(x) - f(y) \le \frac {\epsilon}2 \\ \text { if both } x,y \text { are } in (c, b)\\ \] now let \[ x \in (a, c)\\ y \in (c,b)\\ |x - y | < \delta\\ \text { implies }\\ |x-c|\le \delta\\ |y-c| \le \delta\\ \text{ which implies }\\ |f(x) - f(y)| \le |f(x) - f(c)| + | f(c) - f(y)| \le \le \frac {\epsilon}2 + \le \frac {\epsilon}2=\epsilon \]

OpenStudy (anonymous):

If x and y are in the same sub-interval, there is nothing to prove.